# Marginal Debiased Network for Fair Visual Recognition

### v1

@@ Abstract
publicationid: pubid: $0$–$00$ 2021 IEEE
I Introduction
Recent deep neural networks (DNNs) [1, 2, 3] have shown remarkable performance and have been deployed in some high-stake applications, including criminal justice and loan approval. However, many concerns have arisen that they often show discrimination towards bias attributes like gender and race, leading to serious societal side effects [4, 5, 6, 7]. For example, COMPAS used by the United States judiciary is prone to associate African-American offenders with a high likelihood of recidivism compared with Caucasian [8].
@@ I Introduction
A major driver of model bias is the spurious correlations between target classes and bias attributes that appear in a majority of training data, referred to as dataset bias [9]. In the case of hair length classification shown in Fig. 1(a) and 1(b), for example, hair length is highly correlated to gender attribute, i.e., most people with short hair (i.e., target class) are male (i.e., bias attribute) in a dataset. When DNNs focus more on male images with short hair while ignoring few female images with short hair, unintended correlation contained in the majority of data would mislead models to take gender attribute as a cue for classification, leading to high performance for short-haired males but worst-case errors for short-haired females. Besides, since female training images with short hair are rare, they often fail to describe the true distribution with large intra-class variation, which makes models unable to generalize well on new test samples and thereby leads to inaccurate inferences as well. Throughout the paper, we call data with such spurious correlations and constituting a majority of training data bias-aligned samples (e.g., male with short hair), and the other bias-conflicting samples (female with short hair).
How to get rid of the negative effect of the misleading correlations? One intuitive solution is to enforce models to emphasize bias-conflicting samples and improve their feature representations when training. Existing methods either overweight (oversample) them [10, 11, 12] or augment them by generative models [13, 14]. However, the former cannot address intra-class variation and results in limited improvement of generalization capability, while the latter often suffers from model collapse when training generative models.
In this paper, we focus on the spurious correlations and the generalization gap of different subgroups, and aim to mitigate bias in visual recognition from the perspective of margin penalty. To this end, we propose a marginal debiased network (MDN). Inspired by large margin classification [15, 16], we design a new marginal softmax loss (MSL) which imposes different margins for bias-conflicting and bias-aligned samples. To reduce the negative effects of spurious correlations, our MDN prefers to regularize bias-conflicting samples more strongly than bias-aligned ones so that models would pay more attention to the former and are prevented from learning unintended decision rules from the latter. Furthermore, previous work [17, 18] has demonstrated that margin penalty leads to lower generalization error. Although few bias-conflicting samples are not representative enough to describe the true distribution, margin can reshape their decision boundary and strongly squeeze their intra-class variations to improve generalization on unbiased test criteria.
@@
II Related work
II-A Algorithmic fairness
@@
Bias mitigation techniques have attracted increased attention in recent times. A common practice for mitigating bias is re-weighting or re-sampling [10, 19, 12]. For example, LfF [10] showed that bias-aligned samples are learned faster than bias-conflicting ones, and cast the relative difficulty of each sample into a weight. Ahn et al. [19] claimed that bias-conflicting samples have a higher gradient norm and proposed a gradient-norm-based dataset oversampling method. Another effective way to achieve fairness is by using data augmentation to balance the data distribution [13, 14]. For example, BiaSwap [13] generated more bias-conflicting images via style-transfer for training. Ramaswamy et al. [14] proposed to perturb vectors in the GAN latent space to generate training data that is balanced for each bias attribute. Additionally, there are several other work to resolve dataset bias by confounding or disentangling bias information. LnL [20] minimized the mutual information between feature embedding and bias to predict target labels independent of bias. EnD [21] devised a regularization term with a triplet loss formulation to minimize the entanglement of bias features. Zhang et al. [22] designed controllable shortcut features to replace bias features such that the negative effect of bias can be removed from target features. Orthogonally to these work above, Barbano et al. [23] and we both introduce margin penalty into debiasing, but our algorithm is different from theirs. Barbano et al. [23] proposed to combine a margin-based $\epsilon$-SupInfoNCE loss and a debiasing regularization called FairKL to learn unbiased features. $\epsilon$-SupInfoNCE defines the margin as the minimal distance between positive and negative samples in contrastive learning and aims at increasing it to achieve a better separation between target classes, whereas we define the margin as the distance of the boundary shift and add it to logits to trade off bias-conflicting and bias-aligned samples. Moreover, the margin in $\epsilon$-SupInfoNCE is a predefined hyper-parameter, whereas we learn the margin adaptively via a meta-learning framework.
II-B Large margin classification
The hinge loss is often used to obtain a “max-margin” classifier, most notably in SVMs [24]. With the development of deep learning, there is also a line of work [25, 26, 27] focusing on exploring the benefits of encouraging large margin in the context of deep networks. Elsayed et al. [28] designed a novel loss function based on a first-order approximation of the margin to enhance the model robustness to corrupted labels and adversarial perturbations. Liu et al. [15, 25] incorporated a margin multiplying with the angle between feature and weight to achieve larger angular separability between classes. Different from the multiplicative margin, CosFace [29] and ArcFace [26] further designed an additive cosine and angular margin to minimize the intra-class variation and enlarge the inter-class distance. To address the problem of class imbalance, Adaptiveface [30] is proposed to adaptively adjust the margins for different classes by a regularization term. Kim et al. [31] proposed an adaptive margin function by approximating the image quality with feature norms, which can be utilized to emphasize samples of different difficulties and improve the recognition performance on low-quality data. C-LMCL [32] combined large margin loss with center loss to learn discriminative features for palmprint recognition. Based on triplet loss, Xie et al. [33] assigned an order-insensitive or two order-aware adaptive margin parameters for each expression pair to improve the model generalization. In this paper, we mitigate bias from the perspective of margin penalty. We encourage larger margins for bias-conflicting samples and incorporate a meta learning framework to learn margins adaptively.
II-C Meta learning
@@
Meta-learning [34], or learning-to-learn, has recently emerged as an important direction for developing algorithms for various deep-based applications, including few-shot learning [35, 36], domain generalization [37, 38] and label denoising [39, 40]. In this scenario, an inner-loop phase optimizes the model parameters using training data and an outer-loop phase updates the parameters on meta-validation data such that the model can generalize well to the validation data. MAML [41] used a meta-learner to learn a good initialization for fast adapting a model to a new task. To further simplify MAML, Reptile [42] removed re-initialization for each task and only used first-order gradient information. Zhang et al. [43] proposed a curriculum-based meta-learning method to consider the hardness and quality of tasks and progressively improve the meta-leaner’s generalization. Sun et al. [44] proposed a novel meta-transfer learning to learn to transfer large-scale pre-trained DNN weights for solving few-shot learning task. Li et al. [37] simulated train/test domain shift via the meta-optimization objective to achieve good generalization ability to novel domains. Bui et al. [45] utilized the meta-training procedure to provide a suitable mechanism to self-learn domain-specific representation. MW-Net [46] learned an explicit weighting function for adaptively assigning weights on training samples to address class imbalance and label noise. In this paper, our objective is to learn an unbiased model using biased data. To achieve this, we employ a biased dataset as the training set and utilize an unbiased dataset as the meta-validation set. We take consideration of fairness, and concurrently update both the model and margin parameters through the nested optimization which ensures better generalization on unbiased test sets.
III Methodology
@@
III-A Marginal softmax loss
@@
We give a simple example to explain our idea. Consider the binary classification ($y\in\{0,1\}$ and $b\in\{0,1\}$) and we have a sample $x$ with $b=0$. When $b=0$, let the images with $y=0$ as bias-conflicting samples and the ones with $y=1$ as bias-aligned samples. As shown in Fig. 3(a), the decision boundary of traditional cross-entropy loss is $\eta_{0}-\eta_{1}=0$, which requires $\eta_{0}>\eta_{1}$ to correctly classify $x$ from target class 0 and similarly from target class 1. However, we want to make the classification rigorous for bias-conflicting samples and relaxed for bias-aligned ones in order to shift the model focus from the latter to the former and mitigate the negative effect of dataset bias. To this end, we introduce the margin via a parameter $\{m_{y,b}\}_{y,b=0}^{C,B}$ for each (target, bias) paired group to quantitatively control the decision boundary, and reshape it into $\left(\eta_{0}-m_{0,0}\right)-\left(\eta_{1}-m_{1,0}\right)=0$ for $x$ with $b=0$, where $m_{0,0}>m_{1,0}$ means a larger margin parameter is assigned for bias-conflicting samples. Thus, we force $\eta_{0}-\left(m_{0,0}-m_{1,0}\right)>\eta_{1}$ to correctly classify bias-conflicting samples from class 0, as shown in Fig. 3(b). It is essentially making the decision more stringent than previous, because $\eta_{0}-\left(m_{0,0}-m_{1,0}\right)$ is lower than $\eta_{0}$. Similarly, if we require $\eta_{1}-\left(m_{1,0}-m_{0,0}\right)>\eta_{0}$ to correctly classify bias-aligned samples from class 1, the constraint would be more relaxed since $\eta_{1}-\left(m_{1,0}-m_{0,0}\right)$ is higher than $\eta_{1}$.
Different from open-set face recognition where the margins [29, 26] are inserted between two classes for feature discrimination, the margin here is designed for fairness in closed-set classification and indicates the distance of the boundary shift, i.e, the space between the decision boundaries of training and testing. As shown in Fig. 3(b), bias-conflicting samples are applied a positive boundary shift while bias-aligned samples receive a negative boundary shift. The reshaped decision boundary in training would strongly squeeze the intra-class variations for bias-conflicting samples, leading to improved generalization ability when testing.
@@
III-B Meta margin learning
@@
where $\mathcal{S}_{C}^{y}$ represents the image set containing the under-represented data in the $y$-th target class, and similarly, $\mathcal{S}_{A}^{y}$ is the image set containing the over-represented data in the $y$-th target class. Therefore, $\mathcal{L}_{ME}$ represents the performance skewness of $\{\theta,\phi\left(m\right)\}$ across different samples. Note that we herein assume that $m$ is directly related to the linear classifier $f_{\phi}^{cls}$ via $\phi\left(m\right)$. Therefore, we learn $m^{*}$ based on this assumption: we tune $m$ to make the linear classifier $\phi\left(m\right)$ generalize well, or in other words, minimize $\mathcal{L}_{ME}$, when evaluated on $\mathcal{D}_{meta}$. We formulate a meta equalized loss minimization problem with respect to $m$ as follows,
@@
$\begin{split}m^{t+1}&=m^{t}-\beta\bigtriangledown_{m}\mathcal{L}_{ME}\left(\theta,\phi\left(m\right)\right)\bigg{|}_{\hat{\theta}^{t+1},\hat{\phi}^{t+1},m^{t}}\\ &=m^{t}-\beta\frac{\partial\phi(m)}{\partial m}\bigg{|}_{\hat{\phi}^{t+1},m^{t}}\frac{\partial L_{ME}(\theta,\phi)}{\partial\phi}\bigg{|}_{\hat{\theta}^{t+1},\hat{\phi}^{t+1}}\\ &=m^{t}+\frac{\alpha\beta}{n}\sum_{i=1}^{n}\frac{\partial^{2}L_{MS}^{i}\left(\theta,\phi,m\right)}{\partial\phi\partial m}\bigg{|}_{\theta^{t},\phi^{t},m^{t}}\frac{\partial L_{ME}(\theta,\phi)}{\partial\phi}\bigg{|}_{\hat{\theta}^{t+1},\hat{\phi}^{t+1}},\end{split}$ (7)
where $\beta$ is the descent step size on $m$. Note that the meta gradient update involves a second-order gradient, i.e., $\frac{\partial^{2}L_{MS}^{i}\left(\theta,\phi,m\right)}{\partial\phi\partial m}$, which can be easily implemented through popular deep learning frameworks like Pytorch. Since we make $m$ only directly related to $f_{\phi}^{cls}$, we just need to unroll the gradient graph of $f_{\phi}^{cls}$ and perform backward-on-backward automatic differentiation, which requires a lightweight overhead. Finally, the updated $m^{t+1}$ is employed to ameliorate $\{\theta,\phi\}$,

### v2

@@ Abstract
1 Introduction
Recent deep neural networks (DNNs) [1] have shown remarkable performance and have been deployed in some high-stake applications, including criminal justice and loan approval. However, many concerns have arisen that they often show discrimination towards bias attributes like gender and race, leading to serious societal side effects [2]. For example, COMPAS used by the United States judiciary is prone to associate African-American offenders with a high likelihood of recidivism compared with Caucasian [3].
@@ 1 Introduction
A major driver of model bias is the spurious correlations between target classes and bias attributes that appear in a majority of training data, referred to as dataset bias [4]. In the case of hair length classification shown in Fig. 1(a) and 1(b), for example, hair length is highly correlated to gender attribute, i.e., most people with short hair (i.e., target class) are male (i.e., bias attribute) in a dataset. When DNNs focus more on male images with short hair while ignoring few female images with short hair, unintended correlation contained in the majority of data would mislead models to take gender attribute as a cue for classification, leading to high performance for short-haired males but worst-case errors for short-haired females. Besides, since female training images with short hair are rare, they often fail to describe the true distribution with large intra-class variation, which makes models unable to generalize well on new test samples and thereby leads to inaccurate inferences as well. Throughout the paper, we call data with such spurious correlations and constituting a majority of training data bias-aligned samples (e.g., male with short hair), and the other bias-conflicting samples (e.g., female with short hair).
How to get rid of the negative effect of the misleading correlations? One intuitive solution is to enforce models to emphasize bias-conflicting samples and improve their feature representations when training. Existing methods either overweight (oversample) them [5] or augment them by generative models [6]. However, the former cannot address intra-class variation and results in limited improvement of generalization capability, while the latter often suffers from model collapse when training generative models.
In this paper, we focus on the spurious correlations and the generalization gap of different subgroups, and aim to mitigate bias in visual recognition from the perspective of margin penalty. To this end, we propose a marginal debiased network (MDN). Inspired by large margin classification [7], we design a new marginal softmax loss (MSL) which imposes different margins for bias-conflicting and bias-aligned samples. To reduce the negative effects of spurious correlations, our MDN prefers to regularize bias-conflicting samples more strongly than bias-aligned ones so that models would pay more attention to the former and are prevented from learning unintended decision rules from the latter. Furthermore, previous work [8] has demonstrated that margin penalty leads to lower generalization error. Although few bias-conflicting samples are not representative enough to describe the true distribution, margin can reshape their decision boundary and strongly squeeze their intra-class variations to improve generalization on unbiased test criteria.
@@
2 Related work
2.1 Algorithmic fairness
@@
Bias mitigation techniques have attracted increased attention in recent times. A common practice for mitigating bias is re-weighting or re-sampling [5, 9]. For example, LfF [5] showed that bias-aligned samples are learned faster than bias-conflicting ones, and cast the relative difficulty of each sample into a weight. Ahn et al. [9] claimed that bias-conflicting samples have a higher gradient norm and proposed a gradient-norm-based dataset oversampling method. Another effective way to achieve fairness is by using data augmentation to balance the data distribution [6, 10]. For example, BiaSwap [6] generated more bias-conflicting images via style-transfer for training. Ramaswamy et al. [11] proposed to perturb vectors in the GAN latent space to generate training data that is balanced for each bias attribute. Additionally, there are several other work to resolve dataset bias by confounding or disentangling bias information. Inspired by InfoGAN, LnL [12] proposed to minimize the mutual information between feature embedding and bias, making the features contain less bias information. In practice, it played the minimax game between feature extractor $f$ and bias prediction network $h$. $h$ is trained to predict the bias, while $f$ is trained to make the bias prediction difficult. EnD [13] devised a regularization term with a triplet loss formulation to minimize the entanglement of bias features. Zhang et al. [14] designed controllable shortcut features to replace bias features such that the negative effect of bias can be removed from target features. Orthogonally to these work above, Barbano et al. [15] and we both introduce margin penalty into debiasing, but our algorithm is different from theirs. Barbano et al. [15] proposed a margin-based $\epsilon$-SupInfoNCE loss which defines the margin as the minimal distance between positive and negative samples in contrastive learning and aims at increasing it to achieve a better separation between target classes, whereas we define the margin as the distance of the boundary shift and add it to logits to trade off bias-conflicting and bias-aligned samples. Moreover, the margin in $\epsilon$-SupInfoNCE is a predefined hyper-parameter, whereas we learn the margin adaptively via a meta-learning framework.
@@
2.2 Neural classifier
Recently, neural classifier has been drawing extensive attention. 1) Softmax classifiers learn weight vector and bias term for each class, and have become the de facto regime in computer vision. Our MDN is designed based on Softmax classifier, which helps to achieve fair performance. 2) Centroid-/prototype-based classifiers are nonparametric, classifying data samples based on their proximity to representative prototypes or means of classes. NCM [16] investigated the idea of bringing nearest centroids into CNNs. To discover the latent data structure and handle intra-class variations, DNC [17] further estimated sub-centroids using within-class clustering. Zhou et al. [18] proposed a non-learnable prototype-based semantic segmentation method that represented each class with multiple prototypes and utilized pixel-prototype contrastive learning to achieve better representation. 3) Generative classifiers learn the class densities $p(x|c)$ in contrast to discriminative classifiers which learn the class boundaries $p(c|x)$, where $x$ and $c$ represent data and labels. GMMSeg [19] built Gaussian mixture models via expectation-maximization and learned generative classification with end-to-end discriminative representation in a compact and collaborative manner.
2.3 Large margin classification
The hinge loss is often used to obtain a “max-margin” classifier, most notably in SVMs. With the development of deep learning, there is also a line of work [20] focusing on exploring the benefits of encouraging large margin in the context of deep networks. Elsayed et al. [21] designed a novel loss function based on a first-order approximation of the margin to enhance the model robustness to corrupted labels and adversarial perturbations. CosFace [20] further designed an additive cosine margin to minimize the intra-class variation and enlarge the inter-class distance. Kim et al. [22] proposed an adaptive margin function by approximating the image quality with feature norms, which can be utilized to emphasize samples of different difficulties and improve the recognition performance on low-quality data. Wang [23] presented a novel deep ranking model with feature learning and fusion by learning a large adaptive margin between the intra-class distance and inter-class distance. Our approach is most related to CosFace [20] and Adaptiveface [24] since we both introduce margins into Softmax classifier. However, our objective and algorithm are different from those. 1) CosFace and Adaptiveface aim to learn discriminative features, whereas we focus on achieving fairness. 2) CosFace and Adaptiveface insert the margins between two classes on a hypersphere manifold, while the margins in our MSL indicate the distance of the boundary shift, i.e., the space between the decision boundaries of training and testing. 3) The margins are constants in CosFace, and are adaptively adjusted by a regularization term in Adaptiveface. Different from them, we encourage larger margins for bias-conflicting samples and incorporate a meta learning framework to learn margins adaptively.
2.4 Meta learning
@@
Meta-learning, or learning-to-learn, has recently emerged as an important direction for developing algorithms in various deep-based applications. An inner-loop phase optimizes the model parameters using training data and an outer-loop phase updates the parameters on meta-validation data such that the model can generalize well to the validation data. MAML [25] used a meta-learner to learn a good initialization for fast adapting a model to a new task. To further simplify MAML, Reptile [26] removed re-initialization for each task and only used first-order gradient information. Sun et al. [27] proposed a novel meta-transfer learning to learn to transfer large-scale pre-trained DNN weights for solving few-shot learning task. RMAML [28] developed a geometry aware framework that generalizes the bi-level optimization problem to the Riemannian setting. Li et al. [29] simulated train/test domain shift via the meta-optimization objective to achieve good generalization ability to novel domains. Martins et al. [30] proposed an online tuning of the uncertainty sampling threshold using a meta-learning approach. MW-Net [31] learned an explicit weighting function for adaptively assigning weights on training samples to address class imbalance and label noise. In this paper, our objective is to learn an unbiased model using biased data. To achieve this, we employ a biased dataset as the training set and utilize an unbiased dataset as the meta-validation set. We take consideration of fairness, and concurrently update both the model and margin parameters through the nested optimization which ensures better generalization on unbiased test sets.
3 Methodology
@@
3.1 Marginal softmax loss
@@
We give a simple example to explain our idea. Consider the binary classification ($y\in\{0,1\}$ and $b\in\{0,1\}$) and we have a sample $x$ with $b=0$. When $b=0$, let the images with $y=0$ as bias-conflicting samples and the ones with $y=1$ as bias-aligned samples. As shown in Fig. 3(a), the decision boundary of traditional cross-entropy loss is $\eta_{0}-\eta_{1}=0$, which requires $\eta_{0}>\eta_{1}$ to correctly classify $x$ from target class 0 and similarly from target class 1. However, we want to make the classification rigorous for bias-conflicting samples and relaxed for bias-aligned ones in order to shift the model focus from the latter to the former and mitigate the negative effect of dataset bias. To this end, we introduce the margin via a parameter $\{m_{y,b}\}_{y,b=0}^{C,B}$ for each (target, bias) paired group to quantitatively control the decision boundary, and reshape it into $\left(\eta_{0}-m_{0,0}\right)-\left(\eta_{1}-m_{1,0}\right)=0$ for $x$ with $b=0$, where $m_{0,0}>m_{1,0}$ means a larger margin parameter is assigned for bias-conflicting samples. Thus, we force $\eta_{0}-\left(m_{0,0}-m_{1,0}\right)>\eta_{1}$ to correctly classify bias-conflicting samples from class 0, as shown in Fig. 3(c). It is essentially making the decision more stringent than previous, because $\eta_{0}-\left(m_{0,0}-m_{1,0}\right)$ is lower than $\eta_{0}$. Similarly, if we require $\eta_{1}-\left(m_{1,0}-m_{0,0}\right)>\eta_{0}$ to correctly classify bias-aligned samples from class 1, the constraint would be more relaxed since $\eta_{1}-\left(m_{1,0}-m_{0,0}\right)$ is higher than $\eta_{1}$.
Different from open-set face recognition where the margins [20] are inserted between two classes for feature discrimination, the margin here is designed for fairness in closed-set classification and indicates the distance of the boundary shift, i.e, the space between the decision boundaries of training and testing. As shown in Fig. 3(c), bias-conflicting samples are applied a positive boundary shift while bias-aligned samples receive a negative boundary shift. The reshaped decision boundary in training would strongly squeeze the intra-class variations for bias-conflicting samples, leading to improved generalization ability when testing.
@@
3.2 Meta margin learning
@@
where $\mathcal{S}_{C}^{y}$ represents the image set containing bias-conflicting data in the $y$-th target class, and similarly, $\mathcal{S}_{A}^{y}$ is the image set containing bias-aligned data in the $y$-th target class. Bias-conflicting data refers to the (target, bias) paired group with fewer samples, whereas bias-aligned data is the opposite. Given the labels of target classes and bias attributes, we can easily construct $\mathcal{S}_{C}^{y}$ and $\mathcal{S}_{A}^{y}$ for MEL. Therefore, $\mathcal{L}_{ME}$ represents the performance skewness of $\{\theta,\phi\left(m\right)\}$ across different samples. Note that we herein assume that $m$ is directly related to the linear classifier $f_{\phi}^{cls}$ via $\phi\left(m\right)$. Therefore, we learn $m^{*}$ based on this assumption: we tune $m$ to make the linear classifier $\phi\left(m\right)$ generalize well, or in other words, minimize $\mathcal{L}_{ME}$, when evaluated on $\mathcal{D}_{meta}$. We formulate a meta equalized loss minimization problem with respect to $m$ as follows,
@@
$\begin{split}m^{t+1}&=m^{t}-\beta\bigtriangledown_{m}\mathcal{L}_{ME}\left(\theta,\phi\left(m\right)\right)\bigg{|}_{\hat{\theta}^{t+1},\hat{\phi}^{t+1},m^{t}}\\ &=m^{t}-\beta\frac{\partial\phi(m)}{\partial m}\bigg{|}_{\hat{\phi}^{t+1},m^{t}}\frac{\partial\mathcal{L}_{ME}(\theta,\phi)}{\partial\phi}\bigg{|}_{\hat{\theta}^{t+1},\hat{\phi}^{t+1}}\\ &=m^{t}+\frac{\alpha\beta}{n}\sum_{i=1}^{n}\frac{\partial^{2}\mathcal{L}_{MS}^{i}\left(\theta,\phi,m\right)}{\partial\phi\partial m}\bigg{|}_{\theta^{t},\phi^{t},m^{t}}\frac{\partial\mathcal{L}_{ME}(\theta,\phi)}{\partial\phi}\bigg{|}_{\hat{\theta}^{t+1},\hat{\phi}^{t+1}},\end{split}$ (7)
where $\beta$ is the descent step size on $m$. Note that the meta gradient update involves a second-order gradient, i.e., $\frac{\partial^{2}\mathcal{L}_{MS}^{i}\left(\theta,\phi,m\right)}{\partial\phi\partial m}$, which can be easily implemented through popular deep learning frameworks like Pytorch. Since we make $m$ only directly related to $f_{\phi}^{cls}$, we just need to unroll the gradient graph of $f_{\phi}^{cls}$ and perform backward-on-backward automatic differentiation, which requires a lightweight overhead. Finally, the updated $m^{t+1}$ is employed to ameliorate $\{\theta,\phi\}$,